OpenStudy (anonymous):

Rewrite the rational exponent as a radical expression. (3^2/3)^1/6

OpenStudy (aum):

\(\Huge (3^{\frac 23})^{\frac 16} = 3^{\frac 23 * \frac 16} =3^{\frac 19} = \sqrt [9]{3}\)

OpenStudy (anonymous):

how did you get the denominator as 9

OpenStudy (aum):

\(\Large \frac 23 \times \frac 16 = \frac {2}{18} = \frac 19\)

OpenStudy (mathmate):

@danibear401 Just in case, is the question: \(\Huge (3^{\frac 23})^{\frac 16} \) or \(\Huge ({\frac {3^2}{3}})^{\frac 16} \) If it's the first case, parentheses are required around "2/3" for proper interpretation.

OpenStudy (aum):

^^^ Good point. Parenthesis required around both fractions 2/3 and 1/6 for my first reply to be valid. In the book or online page there may not be a parenthesis because they can write "vertically" such as \(\Large 3^{\frac 23}\). But when you write "horizontally" you have tp group them together using parenthesis: 3^(2/3). By grouping 2/3 using parenthesis, we know 3 is raised to the whole thing (2/3). But if you write 3^2/3 it will be interpreted as \(\Large \frac {3^2}{3}\).

OpenStudy (anonymous):

its the first one and thank you